OpenStudy (anonymous):

The argument of 1-i sqrt{3}/1+i sqrt{3}

OpenStudy (anonymous):

Hey... Try Rationalising the Form to end up with irrational numbs only in the numerator part... Will Ya ?

OpenStudy (anonymous):

pls explain it

OpenStudy (ikram002p):

is it ? \(\Huge \frac{1-i \sqrt{3}}{1+i \sqrt{3}}\)

OpenStudy (anonymous):

Uh.. Rationalising includes changing the sign of the denominator and multiplyin it Up and down/.... (A+iB)/(A−iB) can be written as (A+iB)(A+iB)/(A-iB)(A+iB)

OpenStudy (ikram002p):

just write it into standers formula of complex numbers which is :- z=x+yi Hint multibly by \(\Huge \frac{1+i \sqrt{3}}{1+i \sqrt{3}} \)

OpenStudy (ikram002p):

sry typo , multibly by \(\Huge \frac{1-i \sqrt{3}}{1-i \sqrt{3}}\)

OpenStudy (anonymous):

Yes... This multiplying is what is called as rationalisation

ganeshie8 (ganeshie8):

or simply subtract the arguments

ganeshie8 (ganeshie8):

argument of numerator = -pi/3 argument of denominator = pi/3 so argument of division = -pi/3 - pi/3 = -2pi/3

OpenStudy (anonymous):

use the conjugacy technique. as @ganeshie8 said

OpenStudy (anonymous):

WOW... Thankz @ganeshie8